A manufacturing company that produces a single product has provided the following data concerning its most recent month of operations: Selling price $ 151 Units in beginning inventory 0 Units produced 2,620 Units sold 2,210 Units in ending inventory 410 Variable cost per unit: Direct materials $ 49 Direct labor $ 21 Variable manufacturing overhead $ 12 Variable selling and administrative $ 11 Fixed costs: Fixed manufacturing overhead $ 96,940 Fixed selling and administrative expenses $ 28,730 The total gross margin for the month under absorption costing is:

$17,680

$128,180

$117,380

$70,720
Answer #1

Fixed manufacturing overhead is treated as product under absorption costing

Cost of Goods Sold per unit

= 49+21+12+(96,940/2620) = 119

Total gross margin = (Sales price - Cogs) * sales units

= (151-119) * 2210

= 70,720

Option D